OpenStudy (anonymous):

Look at this linear equation. 3x + 5y = 30 Which of the following equations can be used to solve for y in the linear equation above?

OpenStudy (anonymous):

y=-3/5x+30 y=3x+6 y=-3/5x+6 y=3/5x+30

OpenStudy (anonymous):

@jcpd910

OpenStudy (anonymous):

3x + 5y = 30 Isolate y 5y = 30 - 3x Divide it all b 5 to isolate y further y = 6 - 3/5x

OpenStudy (anonymous):

which is equivalent to y=-3/5x+6

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

No problem